Witch quadratic equation has roots 7 and -2/3
1
Expert's answer
2019-10-29T11:20:28-0400

To solve the problem I used Vieta's formulas.

The roots r1 , r2 of the quadratic polynomial P(x)=ax2+bx+c satisfy r1+r2=-(b/a), r1*r2=c/a

If a=1, then b= -(r1+r2)=-(7-2/3)=-(6+1/3), c=r1*r2=7*(-2/3)=-14/3.

Answer: x2-(6+(1/3))x-14/3=0